OpenStudy (rayman5766):

Which proportion models this problem? On a soccer team, 5 out of 13 players surveyed said they had two or more siblings. The league has 520 players. Which is the best prediction of the number of players in the league that have two or more siblings? A. x/520 =13/5 B. x/520 = 5/13 C. x/5 = 13/520 D. x/13 = 5/520

OpenStudy (rayman5766):

@campbell_st

OpenStudy (campbell_st):

so you are comparing the ratio with the entire population of soccer player 5/13 = unknown/population.

OpenStudy (rayman5766):

cant you just give me a thing to solve like last time

OpenStudy (rayman5766):

I got 105 but thats not one of the answers

OpenStudy (campbell_st):

oops in our question its 5/13 = x/520 you are basically finding an equivalent fraction

OpenStudy (rayman5766):

so is that the answer

OpenStudy (rayman5766):

wait so its b

OpenStudy (campbell_st):

correct